OpenStudy (anonymous):

Joe pays $275 a month for her apartment. If this is 25% of his monthly salary, how much does he make each month?

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@thushananth01

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

Hmmm...

Parth (parthkohli):

Do you know what \(25\%\) is?

OpenStudy (skullpatrol):

Do you know what percent means?

OpenStudy (anonymous):

Be right back... 10 mins.

OpenStudy (anonymous):

and yes i do

OpenStudy (anonymous):

100%:25% so 4:1. Now you can simply multiply 4 by $275 and that will be your answer

OpenStudy (anonymous):

thanks

OpenStudy (skullpatrol):

Why did you just GIVE the answer?

OpenStudy (anonymous):

yep, glad to help!

OpenStudy (anonymous):

275*25/100?

OpenStudy (skullpatrol):

@Hazzar you did not help at all by just giving the answer.

Parth (parthkohli):

One is not *simply* helped by providing them with an answer.

OpenStudy (anonymous):

guys? do i do that ?

OpenStudy (anonymous):

37 help0wanted 0 275*25/100?

OpenStudy (anonymous):

Could you explain?

OpenStudy (skullpatrol):

What do you want explained?

OpenStudy (anonymous):

How to do it. And you're right, it did not help...

Parth (parthkohli):

Okay, Skullpatrol and I will help you. :-)

OpenStudy (anonymous):

I just do not get it,and Id appreciate explanations

OpenStudy (anonymous):

thanks,parth and skull.

Parth (parthkohli):

The first question is: What do you mean by "percent"?

OpenStudy (anonymous):

outof a hundred.

Parth (parthkohli):

And so what is, "25%"?

OpenStudy (anonymous):

25 out of a hundred

Parth (parthkohli):

So if her income is represented by "hundred", then the amount she paid is twenty-five. Do you get what I mean?

OpenStudy (anonymous):

yes...

Parth (parthkohli):

Sure? :-)

OpenStudy (anonymous):

yeah. you mean $25?

Parth (parthkohli):

The answer is not $25!

Parth (parthkohli):

OK, let me get back to the question. What do we want to find according to the question?

OpenStudy (anonymous):

I am sorry; I am not very good att math

Parth (parthkohli):

No, don't worry. Nobody on this planet is :-)

OpenStudy (anonymous):

we want to find hoow much he makes

OpenStudy (anonymous):

Lol.:)

Parth (parthkohli):

Exactly, so we can always call the thing our questions wants \(x\), correct?

OpenStudy (anonymous):

yes,variable

Parth (parthkohli):

That's the word! Yes, so now let's see what the question means. The question says that \($275\) is \(25\%\) of \(x\). Right?

OpenStudy (anonymous):

oh,righttt!

OpenStudy (anonymous):

let me try to set it up as a proportion,

Parth (parthkohli):

You can do that, but you could also set up an equation. That'd make more sense.

OpenStudy (anonymous):

275/x=25/100?

Parth (parthkohli):

But there are two ways to solve the same problem, so go ahead. :-)

Parth (parthkohli):

Yes, that's correct!

OpenStudy (anonymous):

yay! thanks for helping!

OpenStudy (anonymous):

now ill solve...

Parth (parthkohli):

Just to confirm, what'd you get for \(x\)?

OpenStudy (anonymous):

leg me go to my calculator.one sec.

Parth (parthkohli):

You don't need a calculator for that... :-/

OpenStudy (anonymous):

1100?

Parth (parthkohli):

@help0wanted Please show your work along with the answer. Thanks :-)

OpenStudy (skullpatrol):



OpenStudy (anonymous):

okay. 100*275=27500/25=1100

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

is that right?

Parth (parthkohli):

Correct!

OpenStudy (anonymous):

thanks!